[figure: 9.7 cm]

ELECTRICAL AND MAGNETIC PROPERTIES OF YBa_{2}(Cu_{1−x}As_{x})_{3}O_{7−δ} SUPERCONDUCTING COMPOUNDS

I. Ardelean, I.Gr. Deac, V. Pop, I. Stanca^{∗}, Gh. Borodi^{∗∗} and A. Tisan

Babes-Bolyai University, Faculty of Physics, RO 3400 Cluj-Napoca, Romania ∗Department of Physics, University of Oradea, 3700, Oradea ∗∗Institute of Molecular and Isotopic Technologies, 3400, Cluj-Napoca

The influence of partial substitution of Cu by As in the Y Ba_{2}(Cu_{1−x}As_{x})_{3}O_{7−δ} superconducting compounds 0 < x < 0.1 were studied by using X-ray diffraction, electrical resistivity, dc and ac magnetic susceptibility measurements. The Y Ba_{2}(Cu_{1−x}As_{x})_{3}O_{7−δ} samples were prepared by a standard ceramic method. The structure has been analyzed by X-ray diffraction. As doped samples have orthorhombic-type structure. No unreacted phase were observed. Partial substitution of copper by arsenic ions does not suppress the transition from normal to superconducting states for this range of doping. The T_{c} values, determined from dc electrical resistivity and dc and ac magnetic susceptibility measurements, decrease with increasing As concentration, but these still have large value up to 85 K, when x = 0.1.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : V.Pop

Address for correspondence : Babes-Bolyai University, Faculty of Physics, RO-3400 Cluj-Napoca, Romania Tel: ++4064194315 ext: 176, Fax: ++4064191906

Email address : viorel@phys.ubbcluj.ro